OpenStudy (anonymous):

Write an equation of a line in slope intercept form that is parallel to the line x + 4y = 6 and passes through (-8, 5).

OpenStudy (anonymous):

so first put this equation in slope intercept

OpenStudy (hba):

plug in the points

OpenStudy (anonymous):

4y=-x+6 Divide by 4 y=-1/4+3/2

OpenStudy (anonymous):

*-1/4x

OpenStudy (anonymous):

Already done ^^

OpenStudy (anonymous):

haha sorry i just saw the point given